a metal rod of length 20 cm and area of cross section 0.8*10^-4 m^2 is heated at one end for 100 degree celcius and other end at 30 degree celcius , the quantity of heat flow in 5 minute along the rod is 3.3*10^3 joule determine the thermal conductivity

Grade:12th pass

1 Answers

Varali
6 years ago
As you can see that the heat flowing through the rod in 5 mins(300sec) is actually the heat current=Q/t where Q is the heat flow and t is the time... So 3.3*10^5/300=0.011*10^5jouleper sec . Now heat current = thermal conductivity *(T1-T2)*area of cross section/length of rod . Therefore thermal conductivity = heatflow*length of rod/(T1-T2)* Area of cross section . Answer =0.011 *10^5*0.2 m/(100-30)*0.8*10^-4 =0.3928*10^5